A store is selling two types of jelly beans. One kind is $1.20 pound, and the other is $0.90 per pound. If the merchant wants to sell a mixture of $1.11 per pound, how many of each must be used to make 30 pounds?
Answer by Boreal(15235)

mixture of 30 pounds@$1.11 per pound is
30(1.11)=1.2A+0.9B
But A+B=30
A=30-B
substitute
33.3=36-1.2B+0.9B
-2.7=-0.3B
B=9 pounds
A=21 pounds
check
A is 70% of the beans
$1.11 is 70% of the way between $0.90 and $1.20.